A person pushes horizontally with a force of 220 N on a 61 kg crate to move it across a level floor. The coefficient of kinetic friction is 0.23. (a) What is the magnitude of the frictional force? (b) What is the magnitude of the crates acceleration? (a) Numbe 137.49 Unit Unit (b) Number ,83
Answer #1

(a)

friction force Fr = u*N

Fr = u*m*g = 0.23*61*9.8

Fr = 137.49 N

(b)

F - Fr = m*a

220 - 137.49 = 61*a

a = 1.35 m/s^2

answer